OpenStudy (anonymous):

If f(x) = 2x2 + 4, find f(3). 22 16 −0.5 40

jimthompson5910 (jim_thompson5910):

f(x) = 2x^2 + 4 f(3) = 2(3)^2 + 4 f(3) = 2(9)+4 f(3) = 18+4 f(3) = 22

OpenStudy (anonymous):

f(3)=2(2)+4 f(3)=4+4 f(3)=8